OpenStudy (anonymous):

Express in terms of i. -sqrt -496 i have 4i sqrt 31 is this correct or not

OpenStudy (anonymous):

\[-\sqrt{-496}\]

OpenStudy (kropot72):

\[4i \sqrt{31}\]is correct

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

But is that a negative sign in front of your expression?

OpenStudy (anonymous):

yes it is thats what confused me.

OpenStudy (anonymous):

Im assuming it should be -4i sqrt 31

OpenStudy (kropot72):

In that case this is the correct solution. \[-4i \sqrt{31}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)